Session: Feb 2001 (session 1)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

\(Other\) (A. Cattani at Mon Feb 19 17:44:02 MET 2001) Success

Torun (K.Borkowski at Tue Feb 20 12:32:24 MET 2001) Minor failures
BBC8 continued malfunctioning throughout the experiment. Strong RFI.

Effelsberg (A. Kraus at Tue Feb 20 13:31:32 MET 2001) Minor failures
RFI was present. The scans between UT 1049-1104 and UT1203-1230 were lost: we had to fix a minor problem with the receiver (which should not have influenced the experiment). Between UT1313-1415 the FS reported an antenna position error. This was not genuine, it resulted from a communication problem between FS and the antenna control computer.

Westerbork (Tony Foleu at Tue Feb 20 14:33:22 MET 2001) Minor failures
We observed this with the SINGLE DISH setup (RT7) as the tied array bulk delay was being repaired.

Onsala (A. Polatidis at Wed Feb 21 01:47:20 MET 2001) Success
Some RFI present. No other known problems.

Shanghai (Huang Xinyong at Wed Feb 21 07:00:29 MET 2001) Success
Some RFI present. No other known problems.

Urumqi (Liu Xiang at Wed Feb 21 11:02:58 MET 2001) Success

Jodrell_Bank (Alastair Gunn at Wed Feb 21 16:13:01 MET 2001) Minor failures
Telescope had wrong offsets during this experiment but beam size is large in comparison. Possibly a small drop in power. As usual, extreme RFI, especially in VC7. Also see feedback for 21/18 cm session.

Noto (C.Trigilio at Wed Feb 28 15:36:03 MET 2001) Minor failures
First 8 minuts lost due to problems with the elevation engines. No other problems reported.